### v2

# Experimental Evaluation of a Two-Way Intelligibility Protocol for Human-LLM Interaction

Abstract
Our interest is in the design of software systems involving a human-expert interacting–using natural language–with a large language model (LLM) on data analysis tasks. For complex problems, it is possible that LLMs can harness human expertise and creativity to find solutions that were otherwise elusive. On one level, this interaction takes place through multiple turns of prompts from the human and responses from the LLM. Here we investigate a more structured approach based on an abstract protocol described in Baskar et al. (2024) for interaction between agents. The protocol is motivated by a notion of “two-way intelligibility” and is modelled by a pair of communicating finite-state machines. We provide an implementation of the protocol, and provide empirical evidence of using the implementation to mediate interactions between an LLM and a human-agent in two areas of scientific interest (radiology and drug design). We conduct controlled experiments with a human proxy (a database), and uncontrolled experiments with human subjects. The results provide evidence in support of the protocol’s capability of capturing one- and two-way intelligibility in human-LLM interaction; and for the utility of two-way intelligibility in the design of human-machine systems.
1 Introduction
For over four decades, researchers in human–machine systems have recognised the importance of making machine-generated predictions intelligible to humans. Michie Michie (1982) was among the first to note potential mismatches between human and machine representations. He proposed a classification of machine learning (ML) systems into three types Michie (1988): Weak ML systems are concerned only with improving performance, given sample data. Strong ML systems improve performance, but are also required to communicate what it has learned in some human-comprehensible form (Michie assumes this will be symbolic). Ultra-strong ML systems are Strong ML systems that can also teach the human to improve his or her performance.
This framework, intended for human-in-the-loop contexts, focuses on intelligibility, the ability of the system to explain the “what” and “why” within the human’s window of understanding rather than intelligence. The system’s internal model can use any representation, so long as communication is understandable. Michie’s scheme addresses one-way communication from machine to human, without metrics for intelligibility (later addressed in Ai et al. (2021)) or provisions for human-to-machine communication, which is important in collaborative discovery. This categorisation has also recently informed a similar 3-way categorisation for the use of AI tools in scientific discovery Krenn et al. (2022).
Despite growing interest in human–AI interaction Wienrich and Latoschik (2021); Gammelgård-Larsen et al. (2024); Tsiakas and Murray-Rust (2024); Guan (2024), few models prioritise mutual intelligibility. Explainable AI (XAI) often produces explanations without tailoring them to the intended audience, undermining user confidence when advice is opaque and when the system cannot gauge the user’s comprehension. A recent proposal Baskar et al. (2024) addresses this by defining “two-way intelligibility” as a property of communication protocols between agents, modelled as finite-state machines. Messages are tagged with one of four “R’s”: ratification, refutation, revision, or rejection, and intelligibility is assessed from the sequence of these tags. While Baskar et al. (2024) specifies the protocol with proofs of correctness and termination, no implementation or empirical evidence was provided.
Given the capabilities of large language models (LLMs) to produce human-readable predictions and explanations when properly conditioned, they are promising candidates for such protocols. However, determining the right conditioning information is non-trivial. This paper addresses that gap by implementing the Baskar et al. (2024) protocol restricted to interactions between an LLM-based generator-agent and a human tester-agent, providing a partial but practical realisation of two-way intelligibility in action.
2 Related Work
Iterative Prompting: Work has been done in prompting LLMs iteratively, by decomposing the problem into smaller sub-problems, or paths in a decision tree, such as Chain-of-Thought Wei et al. (2023); Sun et al. (2024), Tree-of-Thought Yao et al. (2023), etc. Our work is not a method to iteratively prompt an LLM; rather, it is a method to analyse and interpret the interactions between a Human and an LLM in a multi-turn fashion, using the notion of intelligibility.
Intelligibility: Intelligibility is the ability to be understood or comprehended. Previous work has proposed using concepts in philosophy, psychology, and cognitive science Miller (2018); Hoffman et al. (2022) to gain insight into intelligible systems. Our aim is to measure intelligibility and identify Strong/Ultra Strong ML systems as described by Michie Michie (1988), and see if the characteristics are shared across tasks. Our implementation allows us to explore interactions between not only Human-ML interactions but also interactions between two ML agents. We achieve this by restricting the Agents to Predict and Explain (PEX) as described in Baskar et al. (2024). This constraint enables us to quantify intelligibility using message tags.
3 The Main Aspects of the PXP Protocol
The PXP protocol is an interaction model described in Baskar et al. (2024). It is motivated by the question of when an interaction between agents could be said to be intelligible to either. Restricting attention to specific kinds of agents–called PEX agents–the paper is concerned with when predictions and explanations for a data instance produced by one PEX agent are intelligible to another PEX agent. The agents are modelled as finite-state machines, and “intelligibility” is defined as a property inferrable from the tagged messages exchanged between the finite-state machines.
An agent $a_{n}$ decides the tag of its message to be sent to agent $a_{m}$ based on comparing its prediction $y_{n}$ and explanation $e_{n}$ for a data instance $x$ to the prediction and explanation it received from $a_{m}$ for $x$ ($y_{m}$ and $e_{m}$ respectively). The comparisons define the truth-value of guards for transitions, which in turn define the message-tags. Informally, if $a_{n}$ agrees with $a_{m}$ on both prediction and explanation, it sends a message with a $\mathit{RATIFY}$} tag. If $a_{n}$ disagrees with either the prediction or explanation but not both, then it might change (“revise”) its (internal) model. If it can revise its model, then its message has a $\mathit{REVISE}$} tag; else it has a $\mathit{REFUTE}$} tag. Finally, if $a_{n}$ disagrees with both prediction and explanation, it sends a $\mathit{REJECT}$} tag. In all cases, the complete message has the tag, along with $a_{n}$’s current prediction and explanation. Intelligibility is then defined as properties inferable from observing the message-tags exchanged. The important definitions in Baskar et al. (2024) are:
Definition 1 (One-Way Intelligibility)
Let $S$ be a session between compatible agents $a_{m}$ and $a_{n}$ using $\tt{PXP\,}$}. Let $T_{mn}$ and $T_{nm}$ be the sequences of message-tags sent in a session $S$ from $m$ to $n$ and from $n$ to $m$. We will say $S$ exhibits One-Way Intelligibility for $m$ iff (a) $T_{mn}$ contains at least one element in $\{{\mathit{RATIFY}},{\mathit{REVISE}}\}$; and (b) there is no $\mathit{REJECT}$} in the sequence $T_{mn}$. Similarly for $n$.
Two-way intelligibility follows directly if a session $S$ is one-way intelligible for $m$ and one-way intelligible for $n$. The authors in Baskar et al. (2024) then use these definitions to suggest a Michie-style categorisation of Strong and Ultra-Strong Intelligibility. Based on this suggestion, we define the following:
Definition 2 (Strong and Ultra-Strong Intelligibility)
Let $S$ be a session between compatible agents $a_{m}$ and $a_{n}$. Let $T_{mn}$ and $T_{nm}$ be the sequences of message-tags sent in a session $S$ from $m$ to $n$ and from $n$ to $m$. We will say $S$ exhibits Strong Intelligibility for $m$ iff every element of $T_{mn}$ is from $\{{\mathit{RATIFY}},{\mathit{REVISE}}\}$. We will say $S$ exhibits Ultra-Strong Intelligibility for $m$ iff $S$ exhibits Strong Intelligibility for $m$ and there is at least one element of $T_{mn}$ is a $\mathit{REVISE}$}. Similarly for $a_{n}$.
Similarly for $a_{n}$.We now examine a simple implementation of PXP for modelling interactions in which $a_{m}$ and $a_{n}$ are agents that have access to predictions and explanations originating from LLMs and human-expertise, respectively. For simplicity, we refer to the former as the “machine-agent” and the latter as the “human-agent”, and the interaction between $a_{m}$ and $a_{n}$ as “human–LLM interaction”.
4 Modelling Human-LLM Interaction
The implementation described here is restricted to interaction between a single human-agent and a single machine-agent. The protocol in Baskar et al. (2024) is akin to a plain-old-telephone-system (POTS), and it is sufficient for our purposes to implement it as a rudimentary blackboard system with a simple scheduler that alternates between the two agents. The blackboard consists of 3 tables accessible to the agents. The tables are: (a) $Data$. This is a table consisting of $(s,x)$ pairs where $s$ is a session ID, and $x$ is a data instance; and (b) $Message$. A table of 5-tuples $(s,j,\alpha,\mu,\beta)$ where: $s$ is a session identifier, $j$ is a message-number, $\alpha$ is a sender-id, $\mu$ is a message and $\beta$ is a receiver-id; and (c) $Context$. This a table consisting of 3-tuples $(s,j,c)$ where $s$ is a session-id, $j$ is the message number, and $c$ is some domain-specific context information. For simplicity, message-numbers will be assumed to be from the set $\{0,1,2,\ldots\}$; $\alpha,\beta$ are from $\{h,m\}$ where $h$ denoting “human” and $m$ denotes “machine”; and messages $\mu$ are $(l,y,e)$ tuples, where $l$ is from $\{{\mathit{RATIFY}},{\mathit{REFUTE}},{\mathit{REVISE}},{\mathit{REJECT}}\}$, and $y$ and $e$ are the prediction and explanation respectively. We treat the blackboard as a relational database $\Delta$ consisting of the set of tables $\{D,M,C\}$. For presentability, $\Delta$ is denoted as a “shared” input in the agent-functions below.
The procedure in Alg. 3 implements the interaction. The interaction is initiated by the machine, with its prediction and explanation for a data instance (one-per-session), and proceeds until all data instances have been examined. The function $\mathtt{ASK\_AGENT}$} (Alg. 2) asks the corresponding agent to obtain the prediction and explanation from the corresponding agent. The assessment is a message tag which is obtained using the $\mathtt{AGENT}$} (Alg. 1) procedure. We assume that both agents only send a $\mathit{REJECT}$} tag after the interaction has proceeded for some minimum number of messages. Until then, the machine’s message will be tagged either as $\mathit{REVISE}$} or $\mathit{REFUTE}$}. After the bound, the machine can send a message with a $\mathit{REJECT}$} tag. Similarly, the human-agent will send a $\mathit{REFUTE}$} message to the machine until this bound is reached, after which the message tag can be $\mathit{REJECT}$}. Since the message-length is bounded, it will not affect the termination properties of the bounded version of PXP . The procedure for calling the human- or machine-agent is in Alg. 1. Unsurprisingly, the same procedure suffices for both kinds of agents, since PXP is a symmetric protocol that does not distinguish between agents (other than a special agent called the oracle). Agent-specific details arise in the $\mathtt{MATCH}$} and $\mathtt{AGREE}$} relations, and in the question-answering step (the $\mathtt{ASK\_AGENT}$} function), which is shown in Alg. 2. The $\mathtt{ASSEMBLE\_PROMPT}$} is domain-dependent, and not described algorithmically here. Instead, we present it by example below. The $\mathtt{MATCH}$} and $\mathtt{AGREE}$} functions used are described in section, Sec. 5. We next report results from experiments using this implementation.
Algorithm 1 The $\mathtt{AGENT}$} Procedure.
Input: $q$ an agent query; $\lambda$: an agent identifier; $s$: a session identifier; $j$: a message-number ($j>0$); $k$; message-number after with $\mathit{REJECT}$} tags can be sent $(k>1)$; $\Delta$: a shared relational database;
Output: $(l,y,e)$, where $l\in\{{\mathit{INIT}},{\mathit{RATIFY}},{\mathit{REJECT}},{\mathit{REVISE}},{\mathit{REFUTE}}\}$, $y$ is a prediction, and $e$ is an explanation
1: Let $\Delta=\{D,M,C\}$
2: Let $(s,x)\in D$
[…]
5:if $(j\geq 2)$ then
6:  Let $(s,j-1,\alpha,(l^{\prime},y^{\prime},e^{\prime}),\lambda)\in M$
7:  if $(j=2)$ then
8:   $y^{\prime\prime}:=y$
9:   $e^{\prime\prime}:=e$
10:  else
11:   Let $(s,j-2,\lambda,(l^{\prime\prime},y^{\prime\prime},e^{\prime\prime}),\alpha)\in M$
12:  end if
13:  $CatA:=({\mathtt{MATCH}}_{\lambda}(y^{\prime},y^{\prime\prime})\wedge{\mathtt{AGREE}}_{\lambda}(e^{\prime},e^{\prime\prime}))$
14:  $CatB:=({\mathtt{MATCH}}_{\lambda}(y^{\prime},y^{\prime\prime})\wedge\neg{\mathtt{AGREE}}_{\lambda}(e^{\prime},e^{\prime\prime}))$
15:  $CatC:=(\neg{\mathtt{MATCH}}_{\lambda}(y^{\prime},y^{\prime\prime})\wedge{\mathtt{AGREE}}_{\lambda}(e^{\prime},e^{\prime\prime}))$
16:  $CatD:=(\neg{\mathtt{MATCH}}_{\lambda}(y^{\prime},y^{\prime\prime})\wedge\neg{\mathtt{AGREE}}_{\lambda}(e^{\prime},e^{\prime\prime}))$
17:  $Changed:=(\neg{\mathtt{MATCH}}_{\lambda}(y,y^{\prime\prime})\vee\neg{\mathtt{AGREE}}_{\lambda}(e,e^{\prime\prime}))$
18:  if $(CatA)$ then
19:   $l:={\mathit{RATIFY}}$
20:  else if $(CatB$ or $CatC)$ then
21:   if $(Changed)$ then
22:    $l:={\mathit{REVISE}}$
23:   else
24:    $l:={\mathit{REFUTE}}$
25:   end if
26:  else if $(CatD)$ then
27:   if $(j>k)$ then
28:    $l:={\mathit{REJECT}}$
29:   else if ChangedChanged then
30:    $l:={\mathit{REVISE}}$
31:   else
32:    $l:={\mathit{REFUTE}}$
33:   end if
34:  end if
35:else
36:  $l:=Init$
37:end if
[…]
40:return  $(l,y,e)$
Algorithm 2 The $\mathtt{ASK\_AGENT}$} Function.
Input: $q$: an agent query; $x$: a data instance; $\lambda$: an agent identifier; $C$: prior information;
Output: $(y,e)$, where $y$ is a prediction, and $e$ is an explanation
1:if $(\lambda$ is an LLM)) then
2:  $P:={\mathtt{ASSEMBLE\_PROMPT}}(q,C)$
3:  $(y,e):=\lambda(x|P)$  // ask an LLM
4:else
5:  $(y,e)=\lambda(x|q,C)$  // ask the other agent
6:end if
7:return  $(y,e)$
5 Experimental Evaluation
Algorithm 3 The $\mathtt{INTERACT}$} Procedure.
Input: $X$: a set of data instances; $h$: a human-agent identifier; $m$: a machine-agent identifier; $q_{h}$: a query for the human-agent; $q_{m}$: a query for the machine-agent; $n$: an upper-bound on the total number of messages in an interaction $(n>0)$; $k$: message-number after which an agent can send $\mathit{REJECT}$} tags in messages $(k\geq 1$);
Output: A relational database $\Delta=\{D,M,C\}$
1:$Left:=X$
[…]
3: Let $\Delta=\{D,M,C\}$
4: Share $\Delta$ with $h,m$
5:while ($Left\neq\emptysetdo$
6:  Select $x$ from LeftLeft
7:  Let $s$ be a new session identifier
8:  $D:=D\cup\{(s,x)\}$
9:  $j:=1$
10:  $l_{m}=l_{h}={\mathit{INIT}}$  // dummy assignment
11:  $Done:=(j>n)$
12:  while ($\neg Done$do
13:   $\mu_{m}=(l_{m},y_{m},e_{m}):={\mathtt{AGENT}}(q_{m},m,s,j,k,\Delta)$
14:   $M:=M\cup\{(s,j,m,\mu_{m},h)\}$
15:   $j:=j+1$
16:   $Stop:=((l_{m}={\mathit{RATIFY}}\wedge l_{h}={\mathit{RATIFY}})\vee(l_{m}={\mathit{REJECT}}))$
17:   $Done:=((j>n)\vee Stop)$
18:   if $\neg Done$ then
19:    $\mu_{h}=(l_{h},y_{h},e_{h}):={\mathtt{AGENT}}(q_{h},h,s,j,k,\Delta)$
20:    $M:=M\cup\{(s,j,h,\mu_{h},m)\}$
21:    $j:=j+1$
22:    $Stop:=((l_{m}={\mathit{RATIFY}}\wedge l_{h}={\mathit{RATIFY}})\vee(l_{h}={\mathit{REJECT}}$))
23:    $Done:=((j>n)\vee Stop)$
24:   end if
25:  end while
26:  $Left:=Left\setminus\{x\}$
27:end while
28:return $\Delta$
In this section, we examine Human-LLM interaction through the use of the INTERACT procedure. Experiments reported here look at two real-world problems: X-Ray Diagnosis (RAD), and Molecule Synthesis (DRUG). In the former, we want to use LLMs for diagnosing X-rays and producing reports, and in the latter, we want proposals for synthesis pathways for molecules. We report on two kinds of experiments. First, controlled experiments are conducted using a database of human-authored predictions and explanations and an LLM. This allows us to perform repeated (simulated) experiments, which are needed since sampling variations can arise with the use of LLMs. Secondly, we report on results obtained in uncontrolled experiments using human subjects with varying levels of expertise interacting with an LLM. We use the experiments to assess the following: Human- and Machine-Intelligibility. We estimate the proportion of interactions exhibiting:(a) one- and two-way intelligibility; and(b) Strong- and Ultra-Strong Intelligibility. Machine-Performance. We estimate the changes in machine performance to increase when the interaction is at least one-way intelligible for the agent.111 Human-Performance cannot be assessed with controlled experiments, as the human proxy is a database that is not updated. Human-Performance can be assessed in uncontrolled experiments but it is difficult to establish baselines for comparative changes between specialists in complex domains with differing expertise. The results from uncontrolled experiments allow us to examine evidence to show that the simulation results are representative of real-life usage of the PXP protocol.
5.1 Problems
The RAD problem is concerned with obtaining predictions and explanations for up to 5 diseases from X-ray images. In this paper, an LLM is used to generate diagnoses and reports, given image data. We use data from the Radiopedia database Radiopaedia (2007). The DRUG problem is concerned with obtaining predictions and explanations for the synthesis of small molecules. Once potential molecules (leads) have been identified for inhibiting a target protein, we are interested in synthesising these molecules for testing on biological samples. In this paper, the DRUG problem examines the use of LLMs to propose plans for synthesis, and an explanation of why the plan is proposed. We use data from DrugBank Wishart et al. (2018). Dataset details are provided in Appendix B, and excerpts of sessions demonstrating the protocol’s use in Appendix D.
5.2 Agents
For RAD, the “human-agent” is an agent that has access to the database of the human-derived predictions and explanations. In addition, to obtain the message tags we require the agent includes a $\mathtt{MATCH}$} function for comparing predictions and a $\mathtt{AGREE}$} function for comparing explanations. We implement these as follows: $\mathtt{MATCH}$}: Check whether the prediction of the machine-agent (disease/pathway) matches with the ground truth. This is a simple equality check. $\mathtt{AGREE}$}: Check whether the explanation produced by the machine-agent is consistent with the explanation provided by the human-compiled database. This involves whether the two explanations concur. This is done here by querying a second (“tester”) LLM with a prompt similar to “Are these two reports/pathways consistent with each other?”. We note that since the predictions and explanations are from an immutable database, there is no possibility of the human-agent in RAD sending a $\mathit{REVISE}$} tag. For DRUG, the human-agent has access to a chemist (one of the authors of this paper) who assesses the predictions and explanations from the LLM. We assume the chemist nominally employs $\mathtt{MATCH}$} and $\mathtt{AGREE}$} functions, these are not implemented as computational procedures. The chemist directly provides the message-tag along with predictions and explanations. The chemist’s message can be tagged with any of the tags allowed in PXP . For RAD and DRUG, the machine-agent uses an LLM for generating predictions and explanations, and the machine-agent uses a separate LLM is used to implement $\mathtt{AGREE}$}. The task of this second (“checker”) LLM is to determine if the explanation obtained from the human-agent is consistent with the explanation generated by the machine-agent.
5.3 Method
We consider two kinds of experimental settings. The purpose of controlled experiments is to be able to perform repetitions. These are conducted by way of simulations with a human-proxy, using a database of human-authored predictions and explanations. These experiments are conducted for both RAD and DRUG. We also conduct uncontrolled experiments, using 2 human chemists, with different levels of chemical expertise (consistent with graduate- and doctoral-level training), and with different level of computational expertise. For all experiments, our method below is straightforward:
  1. 1.
    For $r=1\ldots R$:
    1. (a)
      Initiate the INTERACT procedure with the data available obtain a record $\Delta$ on termination of the INTERACT procedure
    2. (b)
      Store $\Delta$ as $\Delta_{r}$
  2. 2.
    Using the records in $\Delta_{1},\ldots,\Delta_{R}$, obtain the frequency of sessions that are: (a) one-way and two-way Intelligible; and (b) Strong and Ultra-Strong Intelligible
  3. 3.
    Estimate the proportions of interest from the median values obtained above
The controlled experiments consist of 5 repetitions with 20 instances (X-ray images for RAD, and small molecules for DRUG). Uncontrolled experimental results are provided with DRUG (radiologists were unavailable). In Baskar et al. (2024), it is assumed that agents employ a LEARN function to decide whether to revise their predictions, based on estimating if revision would improve performance. In controlled experiments, we incorporate this through the use of a validation-set. Thus, the LLM only revises its prediction if its validation-set accuracy improves. In contrast, the human-proxy functions as an oracle (predictions and explanations are always taken to be correct). In uncontrolled experiments, no such set is used by either agent. The INTERACT procedure requires a bound on the total messages exchanged, and is set to $10$. We will use the usual (maximum-likelihood) estimate of proportion as the ratio of the number of sessions with a property to the total number of sessions; We use Claude 3.5 Sonnet as the LLM, with more details in Appendix A.
5.4 Results
We focus first on the controlled experiments. For simplicity, we will continue to call the human-proxy agent as “the human agent”. The statistics of interaction on controlled experiments are tabulated in Tab. 1, Fig. 12. Broadly, they are consistent with the following claims: (a) The proportion of one-way and two-way intelligible sessions for the human agent increases as the length of interaction increases. (b) Machine-performance (measured on test data) increases with the length of interaction. The following additional observations have possibly more interesting long-term consequences of using the $\tt{INTERACT\,}$} implementation of PXP : One- and Two-Way Intelligibility: The frequencies of sessions that are 1-way intelligible are high for both human- and machine-agents in RAD, and very high in DRUG. Some of this is due simply to the vast store of prior data and information within an LLM that allows it to provide correct predictions and explanations almost immediately. For example, in 8 of 20 sessions in DRUG, the chemist and machine-agent agree on the prediction and explanations within 3 exchanges. However, interestingly, in an additional 7 sessions, they agree after some exchange of refutations and revisions. For human-in-the-loop systems, this is indicative of the advantage to the human of being provided with explanations in natural language. It is also evidence that the LLM is able to act on text-based feedback from the human, consistent with the findings in Brown et al. (2020). The number of sessions that are 2-way intelligible clearly cannot be more than the lower number of 1-way intelligible sessions. Strong and Ultra-Strong Intelligibility: While it is not surprising to find the number of strongly intelligible sessions is lower than the number of 1- or 2-way intelligible sessions, the difference in numbers between strong- and ultra-strong sessions is surprisingly high. Resolving this requires examining first the sessions exhibiting strong intelligibility. Closer examination shows that these are exactly those sessions which are immediately ratified by both human- and machine-agents [^2] For all of these the message-tag sequence is: $\langle{\mathit{INIT}}_{m},{\mathit{RATIFY}}_{h},{\mathit{RATIFY}}_{m}\rangle$.. This is a degenerate form of strong-intelligibility (see Defn. 2). More interesting, strong-intelligibility would arise from longer sessions (for example, having a hypothetical sequence of message-tags: $\langle{\mathit{INIT}}_{m},{\mathit{REVISE}}_{h},$ ${\mathit{REVISE}}_{m},{\mathit{RATIFY}}_{h},{\mathit{RATIFY}}_{m}\rangle$). In fact, no such sequences occur. Thus, if we ignore these very short interactions, there are, in fact, very few strongly intelligible sessions. It is interesting though that in DRUG, we are able to observe 18 such sessions for the machine-agent.Variability. The experiments were repeated 5 times. It is also evident from the error bars that sampling variation decreases as the length of sessions increases. We believe this to be due to the increasing context information being available to the generator LLM used by the machine-agent as session-length increases.
Table 1: Interaction statistics. For RAD, the median of 5 runs is reported. Only 1 run was possible for DRUG. Parentheses indicate proportions of the total sessions. In the uncontrolled experiments for DRUG, Human h1 has lower chemical but higher computational expertise than Human h2.
Count RAD DRUG DRUG-h1 DRUG-h2
Total sessions 20 20 20 20
1-way intelligible Human 19 (0.95) 18 (0.90) 19 (0.95) 17 (0.85)
sessions for: Machine 20 (1.00) 20 (1.00) 20 (1.00) 19 (0.95)
2-way intelligible sessions: 19 (0.95) 18 (0.90) 19 (0.95) 17 (0.85)
Strong intelligible Human 4 (0.20) 8 (0.40) 12 (0.60) 7 (0.35)
sessions for: Machine 19 (0.95) 19 (0.95) 20 (1.00) 18 (0.90)
Ultra-Strong intelligible Human 0 (0.00) 0 (0.00) 0 (0.00) 0 (0.00)
sessions for: Machine 15 (0.75) 10 (0.50) 8 (0.40) 11 (0.55)
(a) RAD
(b) DRUG
(c) DRUG - uncontrolled
Figure 1: Human- and Machine-Intelligibility in (a,b) controlled, and (c) uncontrolled experiments. The proportion of one-way intelligible sessions increases as the length of interaction is increased. In RAD, by message 3, 13 sessions are one-way intelligible for the human. Error bars are for 5 repetitions.
(a) Controlled
(b) Uncontrolled
Figure 2: Machine-Performance in (a) controlled, (b) uncontrolled experiments.
We turn now to results from the uncontrolled experiments, shown in Fig. 2(b). Broadly, we see the main trend obtained in controlled experiments repeated here: The proportion of intelligible sessions. It is interesting however, to note that the human-agent with higher computational–not chemical–expertise has a greater proportion of intelligible sessions. Closer inspection showed this to be a chance effect of the LLM’s output (the same molecule can produce very different synthesis plans from the LLM). Thus, while the LLM’s output remains broadly intelligible, the level of intelligibility can vary from one run to another. The Machine-Performance plot also brings out an interesting aspect. In the controlled experiments, the LLM’s prediction was only revised if performance on a validation-set improved. This check is not done in the uncontrolled setting, thus violating one of the assumptions of the protocol in Baskar et al. (2024). The result is that if the assumption is violated, then we can end up in the paradoxical position of the machine finding the human response intelligible, but that not being reflected in improvements in performance.
6 Conclusion
Recently, we have seen a dramatic expansion in the use of ML, enabling it to support almost any activity where data can be collected and analysed. A challenge arises when the predictive power of modern ML meets the need for human understanding. On the face of it, the use of machine-agents that communicate in natural language can mitigate the problem by providing explanations in a human-readable form. But, readability, while necessary for understandability, may not be sufficient. In this paper, we investigate the use of an information-exchange protocol specifically designed around the notion of intelligibility of messages exchanged between agents, that attempts to address the broader issue of the quality of information exchanged in human–machine communication Coiera (2001).
The concept of “intelligibility of communication” underpins the abstract interaction model proposed in Baskar et al. (2024), which explores how human–ML systems can collaborate to predict and explain data. While that work is conceptual, with no implementation, it presents case studies showing how an “intelligibility protocol” could qualitatively assess communication clarity between humans and machines.
This paper builds on that idea by implementing a simple interactive procedure that exchanges messages as proposed in Baskar et al. (2024) and testing it on problems involving human interaction with large language models (LLMs). LLMs are well-suited for this because (1) their natural language capabilities enable effective collaboration with non-ML experts who have domain expertise, and (2) foundational LLMs possess extensive general knowledge useful for complex, data-driven analysis. The results are a first step toward empirical support for using intelligibility as a foundation for designing collaborative human–ML systems.
Appendix A Experiment Details
Here we detail the configuration for the LLMs used in the experiments.
All experiments use claude-3-5-sonnet-latest Anthropic (2024) as the LLM. For ASK_AGENT (Algorithm 2). The max_tokens for RAD is set to 300, and for DRUG it is set to 1024. For AGREE, the temperature parameter for both RAD and DRUG is set to 0, and max_tokens is set to 10. All other settings are the defaults, set by the Claude API.
Appendix B Dataset Details
RAD: We use Radiopaedia  Radiopaedia (2007) for data. This is a multi-modal dataset compiled and peer-reviewed by radiologists. Each tuple in our database consists of: (a) An X-ray image; (b) A prediction consisting of a set of diagnosed diseases, and (c) an explanation in the form of a radiological report. Both (b) and (c) are from expert human annotators. We use a subset of the full Radiopedia database, focusing on 5 ailments ($Atelectasis$, PneumoniaPneumonia, PneumothoraxPneumothorax, PleuralEffusionPleuralEffusion, and $Cardiomegaly$) as our database with 4 instances per disease (20 instances overall). We summarise the reports using gpt-3.5-turbo-0125 and use these as the ground truth.
DRUG: Molecules are selected from DrugBank Wishart et al. (2018) (drugs are leads that satisfy additional biological and commercial constraints). It contains a list of drug molecules approved by the FDA (Food and Drug Administration). Specifically, we will focus on a subset of molecules with a molecular weight between 150g/mol - 300g/mol and have at least one aromatic ring. This set has been further sampled to get 20 molecules.
Appendix C Message Tags
The table below shows how a message is tagged in the AGENT procedure (Algorithm 1), in a simplified format.
Explanation
${\mathtt{AGREE}}(e_{n},e_{m})$ $\neg{\mathtt{AGREE}}(e_{n},e_{m})$
${\mathtt{MATCH}}(y_{n},y_{m})$ $\mathit{RATIFY}$} $\mathit{REFUTE}$} or
Pred. (A) $\mathit{REVISE}$} (B)
$\neg{\mathtt{MATCH}}(y_{n},y_{m})$ $\mathit{REFUTE}$} or $\mathit{REJECT}$}
$\mathit{REVISE}$} (C) (D)
Figure 3: Message tag categories
Appendix D Example Sessions
To illustrate the use of the $\mathtt{INTERACT}$} procedure (Alg. 3), we show excerpts of the conversations betweeen agents for each of the tasks, Fig.4 for RAD and Fig.5 for DRUG. Note that the first input is provided to both the Human and Machine agent at the start with the $\mathit{INIT}$} tag, and is placed between the ‘Human’ and ‘Machine’ columns. The ‘Comments’ column contains the prompts provided.
Figure 4: Excerpt of a session in the RAD experiment
Figure 5: Excerpt of a session in the DRUG experiment